OpenStudy (anonymous):

Complete the statement If x/7 = 5/3, then x+7/7 =?

OpenStudy (anonymous):

x+7/7=12

Parth (parthkohli):

Find the value of x. \(\Large \color{MidnightBlue}{\Rightarrow 3x = 7*5 }\) \(\Large \color{MidnightBlue}{\Rightarrow 3x = 35 }\) \(\Large \color{MidnightBlue}{\Rightarrow x = {35 \over 3} }\) Now, x + 7: \(\Large \color{MidnightBlue}{\Rightarrow {56 \over 7} }\) \(\Large \color{MidnightBlue}{\Rightarrow 8 }\) Let's find the value of (x + 7)/7 \(\Large \color{MidnightBlue}{\Rightarrow {8 \over 7} }\)

OpenStudy (anonymous):

thanks for the help. I know that the answer you gave is correct if you wanted to solve for x+7/7 but I think they wanted to kind of figure out what other letters/ numbers in the original equation will also equal what x+7/7. For instance, they gave an example saying If a/b = 9/4 then 4a = 9b. Although maybe your answer is the correct answer, im not all to sure.

OpenStudy (anonymous):